# LumiSculpt: A Consistency Lighting Control Network for Video Generation

### v1

@@ 3.3.2 Lighting Learning
To obtain diverse appearance data, an naive solution is to use additional video data as regular samples to provide the model with diverse appearances. However, it is challenging to obtain a large amount of diverse data with known lighting conditions, so we opt for the manufacture of regular samples based on generative models. As shown in Figure 3(c), we propose a dual-branch structure, including a training branch and a frozen branch, by introducing a frozen foundational denoising model to provide an appearance reference. During training, the dual branches accept the same textual conditions and noisy latents, obtaining predicted noises $\epsilon_{t}$ and $\epsilon_{t}^{reg}$, respectively. The diverse appearances of the pre-trained model are reflected in $\epsilon_{t}^{reg}$. In this way, we achieve low-cost regular sample manufacturing.
Loss Functions
We train the model to learn the overall distribution of the dataset using a simple denoising loss $\mathcal{L}_{denoise}$, which includes lighting information and appearance information from the training videos, the latter of which we do not require. To remove appearance, a quantitative method (e.g., a loss function) capable of measuring appearance consistency without affecting lighting learning is needed. On the one hand, this loss should be able to quantify the appearance consistency of the latents from two videos, and on the other hand, the loss must be independent of planar location, otherwise it will affect the learning of lighting distribution. Inspired by style transfer (Jing et al., 2020), which also needs to measure style consistency without being disturbed by structured information, we propose a disentanglement loss $\mathcal{L}_{dis}$ to measure appearance consistency. We use AdaIN (Huang & Belongie, 2017) to measure the feature distribution patterns of two videos’ latents, reflecting their appearance feature distribution. Aligning the feature distribution of the training branch-generated video with the frozen branch retains the diverse appearance of the original model. With this approach, the generative model globally aligns with the training video under the drive of $\mathcal{L}_{denoise}$, while squeezing out redundant appearance information under the drive of $\mathcal{L}_{dis}$, retaining the generative capability of the original model. The training objective function can be represented as:
$\mathcal{L}_{dis}$ $=||(\sigma(z^{pred}_{0})-\sigma(z^{reg}_{0})||_{2}+||(\mu(z^{pred}_{0})-\mu(z^{reg}_{0})||_{2},$ (1)
$\mathcal{L}_{denoise}$ $=\mathbb{E}_{z_{1:N},\epsilon,c_{t},t}\left[||\hat{\epsilon}(z^{pred}_{1:N},c_{t},t)-\epsilon||^{2}\right],$
$\mathcal{L}_{total}$ $=\mathcal{L}_{denoise}+\beta\mathcal{L}_{dis},$
where $\sigma(\cdot)$ and $\mu(\cdot)$ denote the computation of standard variance and mean, respectively. $z_{0}^{reg}=z_{t}^{reg}-\epsilon_{reg}$ represents the predicted denoised output of the frozen branch at time step $t$, while $z_{0}^{pred}=z_{t}^{pred}-\epsilon_{pred}$ signifies the predicted denoised output of the controlled branch at time step $t$. $N$ denotes the total number of steps, and $c_{t}$ represents the textual condition. $\beta$ is set to 3.0.
4 Experiments
4.1 Experimental Setup
Methods for comparison  We compare our approach with state-of-the-art text-to-video generation methods Open-Sora (Lab & etc., 2024) and relighting method IC-light (Zhang et al., 2024).
Metrics  We employ a variety of quantitative and qualitative metrics to assess the lighting accuracy, inter-frame coherence, and visual-text similarity of generated videos.
Evaluation dataset  We use 500 different light paths and captions not present in the training dataset as conditions to guide the comparative methods in generating evaluation videos.
Implementation details. In all video generation experiments, we use Open-Sora v1.2.0 (Lab & etc., 2024) with the default network architecture. We set a learning rate of $1\times 10^{-4}$. The input video resolution is $640\times 480\times 29$. The training process for each motion requires approximately $800\sim 1500$ iterations using eight NVIDIA A100. The number of inference steps is set to $T=50$ and the guidance scale is set to $w=7.5$.
Table 1: Quantitative experimental results and ablation study results.The best results are marked as bold and the seconds one are marked by underline.
@@ 4.1 Experimental Setup
Figure 4: Comparison results with state-of-the-art methods IC-Light (Zhang et al., 2024) and Open-Sora (Lab & etc., 2024). The classic horizontal and vertical directions for light movement and two brightness levels are tested to achieve a comprehensive qualitative evaluation.
4.2 Quantitative Evaluations
@@
As shown in Table 1, we use five quantitative metrics for evaluation: (1) Frame-wise CLIP image similarity: we utilize the similarity of frame-wise CLIP (Radford et al., 2021) image embeddings to evaluate the semantic-level video coherence. A higher value indicates greater inter-frame similarity, suggesting better semantic stability in the generated video. (2) Frame-wise Learned Perceptual Image Patch Similarity (LPIPS) consistency: we measure feature-level coherence using frame-wise LPIPS consistency. A lower value signifies smaller feature discrepancies, indicating higher inter-frame consistency. (3) Light directions Root Mean Squared Error (RMSE): we calculate the lighting direction for each frame and then assess the consistency of the generated video’s lighting direction with the reference. RMSE represents the discrepancy in lighting direction; the smaller the value, the more consistent the lighting with the target. (4) Brightness consistency: we segment each video frame into patches, compute the average brightness of different patches, and construct a brightness distribution relationship. This distribution is only related to the comparison between different patches and is independent of the absolute brightness values. We calculate the brightness distribution consistency between the generated video and the reference video. (5) CLIP text-image similarity: we measure the model generation quality using the similarity between the clip image embedding of the generated video frames and the text embedding of the caption. The higher the similarity, the better the generation quality. It can be observed that, compared to Open-Sora and IC-Light, LumiSculpt is capable of maintaining good inter-frame consistency and text-image consistency while achieving accurate lighting control.
4.3 Qualitative Evaluations
@@
Figure 5: Ablation study results. We conducted ablations on key components related to model training, including caption augmentation and $\mathcal{L}_{dis}$, both of which contribute to the diversity of generated content. We presented results with varying values of the important parameter guidance scale, which affects the accuracy of lighting and the diversity of appearance, related to model inference.
Table 2: Experimental results of hyper-parameter guidance scale. The best results are marked as bold and the seconds one are marked by underline.
@@ 4.4 Ablation Study
5 Conclusion

### v2

@@ 4.2. Lighting Learning
A straightforward approach to obtaining diverse appearance data is to use additional video datasets as regular samples, thereby exposing the model to a wider range of appearances. However, acquiring large-scale, diverse data with known lighting conditions remains challenging. To address this, we generate regular samples using generative models.
As illustrated in Figure 5(c), we propose a dual-branch framework consisting of a training branch and a frozen reference branch. The frozen branch utilizes a pre-trained foundational denoising model to serve as an appearance reference. During training, both branches receive the same textual condition and noisy latent input, producing predicted noise values $\epsilon_{t}$ and $\epsilon_{t}^{reg}$, respectively. The diverse appearance priors from the pre-trained model are captured in $\epsilon_{t}^{reg}$, enabling the generation of regular samples in a cost-efficient manner.
Table 2. Quantitative experimental results and ablation study results.The best results are marked as bold and the seconds one are marked by underline.
@@
Loss Functions
We train the model to capture the overall distribution of the dataset using a denoising loss $\mathcal{L}_{\text{denoise}}$, which captures both lighting and appearance information. However, since appearance variation is not required, we propose a disentanglement loss, $\mathcal{L}_{\text{dis}}$, to evaluate and suppress appearance information. $\mathcal{L}_{\text{dis}}$ satisfy two criteria: (1) it should effectively quantify the consistency of appearance between the latent representations of two videos, and (2) it should be invariant to planar spatial information, as dependence on pixel location could hinder the learning of lighting distributions. Inspired by neural style transfer (Jing et al., 2020), which similarly aims to preserve style while discarding structural information, we adopt Adaptive Instance Normalization (AdaIN) (Huang and Belongie, 2017) to compare the feature statistics—mean and variance—of the latents from two branches. This effectively captures the distribution of appearance features. By aligning the appearance distribution of the training branch with that of a frozen reference branch, we retain the diverse generative capacity of the original model while filtering out redundant appearance signals. Under this framework, the generative model is optimized to match the training video via $\mathcal{L}_{\text{denoise}}$, while $\mathcal{L}_{\text{dis}}$ enforces appearance disentanglement. Together, they enable precise lighting control while preserving generalization. The overall training objective is defined as:
(1) $\mathcal{L}_{\text{dis}}$ $=\left\|\sigma(z^{\text{pred}}_{0})-\sigma(z^{\text{reg}}_{0})\right\|_{2}+\left\|\mu(z^{\text{pred}}_{0})-\mu(z^{\text{reg}}_{0})\right\|_{2},$
$\mathcal{L}_{\text{denoise}}$ $=\mathbb{E}_{z_{1:N},\epsilon,c_{t},t}\left[\left\|\hat{\epsilon}(z^{\text{pred}}_{1:N},c_{t},t)-\epsilon\right\|^{2}\right],$
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{denoise}}+\beta\mathcal{L}_{\text{dis}},$
where $\sigma(\cdot)$ and $\mu(\cdot)$ denote the standard deviation and mean, respectively. $z_{0}^{\text{reg}}=z_{t}^{\text{reg}}-\epsilon_{\text{reg}}$ and $z_{0}^{\text{pred}}=z_{t}^{\text{pred}}-\epsilon_{\text{pred}}$ represent the predicted denoised latents of the frozen and training branches at timestep $t$. NN is the total number of denoising steps, $c_{t}$ is the textual condition, and $\beta$ is a balancing coefficient set to 3.0.
@@
Figure 6. Comparison results with state-of-the-art methods IC-Light (Zhang et al., 2024), ControlNet (Zhang et al., 2023a) and Open-Sora (Lab and etc., 2024). The classic horizontal and vertical directions for light movement and two brightness levels are tested to achieve a comprehensive qualitative evaluation.
5. Experiments
5.1. Experimental Setup
Methods for comparison
We compare our approach with state-of-the-art text-to-video generation methods Open-Sora (Lab and etc., 2024), image relighting method IC-light (Zhang et al., 2024), and image control method ControlNet (Zhang et al., 2023a).
Metrics
We employ a variety of quantitative and qualitative metrics to assess the lighting accuracy, inter-frame coherence, and visual-text similarity of generated videos.
Evaluation dataset
We use 500 different light paths and captions not present in the training dataset as conditions to guide the comparative methods in generating evaluation videos.
Implementation details.
In all video generation experiments, we use Open-Sora v1.2.0 (Lab and etc., 2024) with the default network architecture. We set a learning rate of $1\times 10^{-4}$. The input video resolution is $640\times 480\times 29$. The training process for each motion requires approximately $800\sim 1500$ iterations using eight NVIDIA A100. The number of inference steps is set to $T=50$ and the guidance scale is set to $w=7.5$.
5.2. Quantitative Evaluations
@@
Table 2 presents five quantitative metrics used for evaluation: (1) Frame-wise CLIP image similarity assesses semantic-level video coherence by measuring the similarity of frame-wise CLIP (Radford et al., 2021) image embeddings. A higher value indicates stronger inter-frame similarity and better semantic stability in the generated video. (2) Frame-wise LPIPS consistency evaluates feature-level coherence by calculating frame-wise LPIPS. A lower value reflects smaller feature discrepancies, indicating higher inter-frame consistency. (3) Lighting direction RMSE is computed for each frame and assess the consistency of the generated video’s lighting direction with the reference. A smaller RMSE indicates better alignment with the target lighting direction. (4) Brightness consistency involves segmenting each video frame into patches and computing the average brightness for each patch to construct a brightness distribution. This distribution is used to measure the consistency between the generated video and the reference, independent of absolute brightness values. (5) CLIP text-image similarity measures the alignment between the generated video frames’ CLIP image embeddings and the text embedding of the caption. A higher similarity indicates better generation quality. As shown in the results, LumiSculpt outperforms Open-Sora-Plan (Lab and etc., 2024) and IC-Light (Zhang et al., 2024) by maintaining strong inter-frame consistency and text-image alignment, while also achieving precise lighting control.
5.3. Qualitative Evaluations
@@
Figure 7. Ablation study results. We conducted ablations on key components related to model training, including caption augmentation and $\mathcal{L}_{dis}$, both of which contribute to the diversity of generated content. We presented results with varying values of the important parameter guidance scale, which affects the accuracy of lighting and the diversity of appearance, related to model inference.
Table 3. Experimental results of hyper-parameter guidance scale. The best results are marked as bold and the seconds one are marked by underline
@@ 5.4. Ablation Study
Generalization
As shown in the Figure 8, LumiSculpt provides lighting priors on animals, which demonstrates the generalization ability to real-world cases.
Figure 8. The results of LumiSculpt on animals.
6. Conclusion